# TempoMaster: Efficient Long Video Generation via Next-Frame-Rate Prediction

### v1

@@ Abstract
We present TempoMaster, a novel framework that formulates long video generation as next-frame-rate prediction. Specifically, we first generate a low-frame-rate clip that serves as a coarse blueprint of the entire video sequence, and then progressively increase the frame rate to refine visual details and motion continuity. During generation, TempoMaster employs bidirectional attention within each frame-rate level while performing autoregression across frame rates, thus achieving long-range temporal coherence while enabling efficient and parallel synthesis. Extensive experiments demonstrate that TempoMaster establishes a new state-of-the-art in long video generation, excelling in both visual and temporal quality.
@@ 3.3 Training
However, during the training phase, we observe that directly employing multi-mask training on multi-shot videos produced unsatisfactory performance. We attribute this failure to information leakage: the random frame sampling in multi-mask training allows the model to observe all shots and thereby weakens the need to learn shot transitions. Thus, for training on multi-shot videos, we randomly select shots and discard all conditioning frames within those shots. This augmentation forces the model to learn to generate new shots, thereby building an inherent capability for multi-shot video generation.

### v2

@@ Abstract
We present TempoMaster, a novel framework that formulates long video generation as next-frame-rate prediction. Specifically, we first generate a low-frame-rate clip that serves as a coarse blueprint of the entire video sequence, and then progressively increase the frame rate to refine visual details and motion continuity. During generation, TempoMaster employs bidirectional attention within each frame-rate level while performing autoregression across frame rates, thus achieving long-range temporal coherence while enabling efficient and parallel synthesis. Extensive experiments demonstrate that TempoMaster establishes a new state-of-the-art in long video generation, excelling in both visual and temporal quality. See our project page at https://scottykma.github.io/tempomaster-gitpage/
@@ 3.3 Training
However, during the training phase, we observed that directly employing multi-mask training on multi-shot videos produced unsatisfactory performance. We attribute this failure to information leakage: the random frame sampling in multi-mask training allows the model to observe all shots and thereby weakens the need to learn shot transitions. Thus, for training on multi-shot videos, we randomly select shots and discard all conditioning frames within those shots. This augmentation forces the model to learn to generate new shots, thereby building an inherent capability for multi-shot video generation.
@@ 5 Conclusion
Appendix
Training Stage Training Steps Batch Size Warmup Steps Weight Decay EMA Weight
Single-Frame-Rate 15000 32 2000 1e-4 0.999
Multi-Frame-Rate 45000 32 2000 1e-4 0.999
Table 5: Additional hyperparameters in training. We enable ema for a more consistent training.
Appendix A Training Details
Base Model
We employ the high-noise model of Wan2.2 [wan2025] as our base model, which can only denoise at an extremely high noise level. In contrast to its predecessor Wan2.1, this model eliminates the dependency on CLIP [clip2021] features extracted from the first frame. This structural simplicity makes it more suitable for our Multi-Mask conditioning framework. Since the base model is inherently limited to extreme noise levels, we first adapt it through a training phase that enables denoising across all noise levels, as mentioned in this paper.
Hyperparameters
We employ the AdamW optimizer with a consistent learning rate schedule, using learning rates of 5e-4 and 2e-5 for the respective training stages. We further provide additional hyperparameters during training in Tab. 5.
Noise Scheduling
During training, we adopt the logit-normal distribution over $t$ following [sd2024]:
$\pi_{\text{ln}}(t;m,s)=\dfrac{1}{s\sqrt{2\pi}}\dfrac{1}{t(1-t)}\exp{-\dfrac{(\text{logit}(t)-m)^{2}}{2s^{2}}}$ (7)
Specifically, we set $m$ to 0 and $s$ to 1, and employ a sigma shift of 3 during training.
Figure 8: Qualitative comparisons. We compare TempoMaster with representative long video methods.
Appendix B Training Data
Our data pipeline draws upon existing short video methods [moviegen2024, seedance2025, wan2025, waver2025]. To ensure the high quality of training data, we filter videos through multi-dimensional quantitative evaluation:
  • Aesthetic Assessment: Video frames are evaluated through the average score from an image aesthetic model. This process guarantees that the training data possesses high aesthetic value, thereby enhancing the visual appeal of the generated content.
  • Clarity Detection: This module primarily employs the Laplacian operator to quantify image sharpness. It focuses on monitoring detail fidelity in dynamic scenes, effectively identifying and filtering out blurred sequences to maintain high image quality across the training samples.
  • Motion Analysis: The coherence and magnitude of motion are assessed by computing optical flow between consecutive frames. This allows for the effective exclusion of static frames and motion-distorted content, thereby improving the dynamic expressiveness of the dataset.
Appendix C The Principle of Human Study
The human evaluation study is designed to assess the quality of generated long videos across multiple key dimensions, reflecting both aesthetic and functional performance. Each video is rated on a scale from 1 to 5 across the following four dimensions:
  • Aesthetic Quality: Evaluates visual appeal through composition, clarity, lighting, and detail rendering. Penalties apply for overexposure, clutter, artifacts, or blurriness. Severe flaws receive the lowest score.
  • Semantic Alignment: Measures fidelity to the text prompt, including background, action, and lighting. Minor deviations result in a 1-point deduction, while partial or complete failure to execute the described actions results in scores of 1–2.
  • Motion Quality: Assesses movement amplitude, speed, and plausibility. Deductions occur for static frames, incoherent motion, implausible dynamics, or severe artifacts.
  • Content Consistency: Tracks temporal degradation in subject appearance, motion degradation (e.g., incoherence or repetition), and visual decay (e.g., color shift or blurring). Significant drift or decay leads to a score of 1.
Appendix D Additional Visualization Results
We provide more qualitative comparison results with prior works in Fig. 8. We observe that FramePack outputs are often characterized by a lack of dynamic motion, whereas the other compared methods exhibit marked degradation in visual quality.